0 0
Add a comment Improve this question Transcribed image text
Answer #1

4) form the conservation of energy

k1+u1 = k2+u2

here u1 = 0

mgh = 1/2m(v1^2-v2^2)

h = 1/2*9.8(33.5^2-31.8^2)

h = 5.66 m

5) Ei = k1+u1 = 1/2*44*140^2+44*9.8*114 = 431200+49156.8 = 480356.8 J

b) wf = Ef-Ei

Ef = 1/2*44*99.2^2+44*9.8*318 = 216494.08+137121.6 = 353615.68 J

wf = -126741.12 J

c) w tot = w+1.5 w = 2.5 wf = -3.17*10^5 J

w tot = k2+u2-k1-k1

w tot = 1/2m(v2^2-v1^2)-mgh1

v2 = sqrt(2wtot+mv1^2+2mgh1/m) = sqrt(-6.34*10^5+8.624*10^5+9.83*10^5/44)

v2 = 165.93 m/s